OpenStudy (anonymous):

SShow the steps and work..Please help <1 and <2 are supplementary angles. m<1 is 4y+7, and m<2 is 9y+4. Find m<2?

OpenStudy (anonymous):

Supplementary means that they add to make 180. So we know that angle 1 + angle 2 = 180 (4y+7) + (9y+4) = 180 Combine like terms 13y + 11 = 180 -11 -11 isolate the variable 13y = 169 y = 13 now substitute back into the equation for angle 2 9(13) + 4 = 121